A person is trying to ride a bike all the way round the inside of



a pipe for a stunt in a film. The filmmaker wants to know what



speeds are involved. The pipe has a diameter of 8 m.



The mass of the bike and rider is 400 kg. The rider goes at a



constant speed of 5 m/s.



a) What is its acceleration at the bottom?



b) What is the force on the bike at an angle of 30° up from the



bottom?



c) What is the minimum velocity at the top for the bike and



rider to stay moving in a circle?



d) Do the bike and rider have sufficient velocity to stay moving



on a circle at the top?�

1
Expert's answer
2022-01-21T08:25:25-0500

(a) "a=g+2v^2/d=16.05\text{ m/s}^2."

(b) "F=m\bigg(g\cos\theta+\dfrac {2v^2}d\bigg)=5895\text{ N}."


(c) "u=\sqrt{gd/2}=6.26\text{ m/s}."

(d) No, because the chosen velocity (5 m/s) is smaller than the required velocity (6.23 m/s).